# Automated Modeling Method for Pathloss Model Discovery

### v3

@@ III-B The Indoor and Outdoor Empirical Models
El Chall et al.[ElChall] derived empirical PL models for LoRaWAN (Long-Range Wide Area Network) based on extensive field measurements conducted in Lebanon. Indoor experiments were carried out at Saint Joseph University, while outdoor experiments were performed in Beirut City. The collected data were analyzed using linear regression to derive PL models for both indoor and outdoor environments.
@@
where $n=2.85$ is the path loss exponent, $PL_{0}=120.4$ is the reference PL, $b=0.47$ is a fitting parameter, $L_{f}=10$ and $L_{w}=1.41$ are the attenuation factors for floors and walls, respectively.
@@ IV Methodologies
The design decision to adopt KANs and DSR as candidate techniques to enable the automation of PL learning is justified by the shortcomings of typical techniques in learning logarithmic functions [makke2024interpretable]. Knowing that existing PL models often involve logarithmic functions in their expression [Phillips], KANs and DSR may have superior performance in approximating such models while being able to produce different flavors of interpretability.
Overall, different methods provide varying levels of interpretability. DNNs offer low interpretability through their activation maps. In contrast, KANs offer a medium level of interpretability through their spline-based representations. DSR and KANs (by using Auto-symbolic mapping) can achieve high interpretability by expressing models as explicit mathematical functions.
@@ IV-A Kolmogorov-Arnold Networks
KANs have recently emerged as an alternative to multi-layer perceptron (MLP) based architectures, demonstrating strong performance on scientific tasks while maintaining interpretability. KANs have strong mathematical foundations based on the Kolmogorov-Arnold representation theorem, which states that any multivariate continuous function can be expressed as a sum of univariate functions [Liukan]. KANs and MLPs implement activation functions in fundamentally different ways: KANs have activation functions on edges, while MLPs have activation functions on nodes. This unique design allows KANs to provide a more transparent and interpretable mapping between inputs and outputs, enhancing explainability while preserving the model’s ability to capture complex patterns.
@@
IV-B Deep Symbolic Regression
At the core of DSR [Petersen] are the representation of mathematical expressions as sequences, the development of an autoregressive model to generate these expressions, and the use of a risk-seeking policy gradient approach to train the model to produce more accurate and well-fitting expressions.
@@ IV-B1 Risk-Seeking Policy Gradient (RSPG)
Originally proposed in [Petersen], this approach prioritizes optimizing the best-case reward rather than the average. The loss function is defined as:
@@ IV-B2 Vanilla Policy Gradient (VPG)
This method employs the REINFORCE algorithm [Williams], where training is performed over the batch $\mathcal{T}$ with the loss function:
@@ IV-B3 Priority Queue Training (PQT)
Proposed by Abolafia [Abolafia], this non-RL approach focuses on optimizing best-case performance. It stores samples from each batch in a maximum reward priority queue (MRPQ), and training is conducted on these stored samples using a supervised learning objective:
@@ V-A1 Synthetic Data
The dataset was generated following the approach described in [anaqreh] to approximate the PL models defined by Eq. (1) and Eq. (2) in Section III-A. The parameter ranges, summarized in Table I, were derived from authentic data in studies that comprehensively analyze various scenarios and frequency ranges, highlighting the applicability of $PL^{ABG}$ and $PL^{CI}$ models in various contexts [Sun, Sun2, Cheng]. Specifically, for $PL^{ABG}$ the parameters $\alpha$$\gamma$, $\beta$, $f$, $d$, and $\chi^{ABG}_{\sigma}$ were used as inputs, while for $PL^{CI}$ $f$,$n$, $d$, and $\chi^{CI}_{\sigma}$ were the inputs. The dataset consists of 1,000 instances. The frequency was converted to hertz for the $PL^{CI}$, following standard practice in the literature.
@@ V-A2 Empirically Measured Data
El Chall et al. [ElChall] conducted measurements in both indoor and outdoor environments. The indoor dataset includes 1,317 instances, and the outdoor dataset contains 787 instances. Input parameters for $PL^{EI}$ included $n_{w}$, $n_{f}$, $d$, and $f$ , while the $PL^{EO}$ used $h_{\mathrm{ED}}$, $d$, and $f$. The corresponding parameter ranges are presented in Table II. For performance evaluation, we compared the automated methods with the models described in Section III-B and developed by the authors in [ElChall] using the real-world measurements [^2] https://zenodo.org/records/1560654.
@@ V-A3 Analytical Baselines
Figure 2: The flow diagram of developing a pathloss model using KANs.
@@ V-B Training of KANs
We used the original KAN implementation from the pykan library [^3] https://github.com/KindXiaoming/pykan and trained the KANs in a similar way as other ML models, using the 80-20 split of data as depicted in Figure 2. During training, layers and grid sizes are configured to optimize performance. In this process, the splines and weights of the edges connecting these splines, as described in Section IV-A, are learned. Unlike MLPs, KANs are not very deep (i.e., usually about 3 layers with 5 being very deep) and not very wide [Liukan]. Due to their relative shallowness and narrowness, once learned, they can be visualized as a graph with the learned activation functions overlayed on the edges and the importance of compositions represented as edge weight. This alone already gives more intuition on the learned model than in the case of other neural architectures. However, as a second step, a mechanism to match the learned splines with actual known functions also exists. We perform default automated matching relying on the library’s default auto_symbolic()auto\_symbolic() functionality. After the matching step, the network is retrained with the respective functions by only updating the weights of the edges. Finally, a fully interpretable symbolic formula is extracted as a last step in the training. To tune the KANs, we perform a full grid search for each model with $grids\in\{5,8,10,15,20,30,40,50\}$, $steps\in\{50,100,200,300\}$ and $lambda\in\{0.02,0.002,0.0002,0.00002\}$.
@@
The final tuned hyperparameters for the PL models learned by KANs are summarized in Table III. For both the CI and ABG models, we normalize the training data to maintain consistency with our previous work [anaqreh]. In preliminary experiments without normalization, the frequency, expressed in hertz in the CI model, tended to dominate the other input variables, resulting in imbalanced feature contributions.
Figure 3: The flow diagram of developing a pathloss model using DSR.
@@ V-C Training of DSR
In DSR, a set of constraints can be implemented to narrow the search space, ensuring the generation of meaningful expressions while preserving computational efficiency. The following constraints were specifically applied. Expression Length Limits: Expressions were constrained to a minimum length of 4 tokens to avoid trivial solutions and a maximum length of 40 tokens to maintain interpretability. Avoiding Redundant Constants: Operators were restricted to prevent all their children from being constants, preventing the generation of expressions that simplify to a single constant. Unary Operator Consistency: Unary operators were restricted from having their inverse as a child (e.g., $\log(\exp(x))$) to prevent the generation of redundant expressions. Trigonometric Operator Composition: Trigonometric operators were restricted from having other trigonometric operators as descendants (e.g., $\sin(1+\cos(x))$) to prevent unnecessary complexity, despite being mathematically valid. Repeat Constraint: This constraint regulates the occurrence of specific tokens (e.g., operators and/or variables) in generated expressions by enforcing minimum and maximum limits. It softly penalizes violations of the minimum threshold to allow exploration while strictly invalidating expressions that exceed the maximum limit. The repeat constraint with $f$ and $d$ was utilized in all cases.
Each of the three RNN-based algorithms (RSPG, PQT, and VPG) has one unique algorithm-specific parameter, for which we adopted default values as presented in [Petersen], i.e., the risk factor ($\epsilon=0.05$) for RSPG, the priority queue size ($k=10$) for PQT, and the EWMA coefficient ($\alpha=0.25$) for VPG. In addition to these unique parameters, all three algorithms share a common set of hyperparameters (batch size, learning rate, and entropy weight). We explored various values for these hyperparameters to achieve the best performance. The final tuned hyperparameters are summarized in Table IV.
@@ VI-A Learning the ABG Model from Synthetic Data
The performance comparison for the KAN and DSR methods on $PL^{ABG}$ is presented in Table V. The KAN [6,6,1] achieves superior predictive performance, with exceptionally low error metrics (MAE = 0.06, MSE = 0.01, MAPE = 0.28) and $R^{2}$ value of 1, indicating a nearly perfect alignment between the model predictions and the actual data as confirmed by the plot in Figure 4. In the figure, pred stands for the predicted PL values and true stands for the actual values based on Eq. (1). In contrast, the KAN [6,6,1] Auto-symbolic, which replaces the learned splines with known functions and re-learns new weights for the edges of the graph, shows significantly lower accuracy, with MAE = 3.89, MSE = 28.14, MAPE = 12.51, and $R^{2}$ = 0.93. Among DSR models, DSR-RSPG obtained the best performance (MAE = 3.86, MSE = 33.50, MAPE = 14.28, $R^{2}$ = 0.92), followed by DSR-PQT (MAE = 5.90, MSE = 62.07, MAPE = 21.83, $R^{2}$ = 0.90) and DSR-VPG (MAE = 6.67, MSE = 83.38, MAPE = 20.48, $R^{2}$ = 0.77). These results highlight that while KAN [6,6,1] achieves the highest accuracy, other methods show varying performance metrics, generally with higher error rates and lower values of $R^{2}$.
@@
TABLE V: Performance comparison on ABG model.
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
KAN [6,6,1]* 0.06 0.01 0.28 1 Tree and splines in Fig. 5 Medium ?
KAN [6,6,1] Auto-symbolic* 3.89 28.14 12.51 0.93 $-291.5\operatorname{log}_{10}{\left(8.7-3.1\alpha\right)}+32.2\cos{\left(1.1\gamma-1.3\right)}+9\beta-53.3\operatorname{log}_{10}{\left(9.4-3.9d\right)}+0.02\left(0.03f-1\right)^{2}+313.3-\frac{0.5}{\left(-f-0.2\right)^{2}}+8\chi_{\sigma}$ High
DSR-RSPG 3.86 33.50 14.28 0.92 $20\alpha+10\gamma+\beta+\log_{10}(d)+\frac{\gamma f}{10}+\chi_{\sigma}$ High
DSR-PQT 5.90 62.07 21.83 0.90 $20\alpha+15\gamma+\beta+\frac{f}{20}+\chi_{\sigma}$ High
DSR-VPG 6.67 83.38 20.48 0.77 $(\alpha+\gamma)(10+\chi_{\sigma})$ High
@@
We examine the validity of the expressions generated by the KANs and DSR methods. For $PL^{ABG}$, the expressions obtained by KAN [6,6,1] Auto-symbolic and DSR-RSPG are considered valid, as they correctly represent PL as a function of $f$ and $d$. We also notice that the expression learned by the KAN [6,6,1] Auto-symbolic is more complex than the one learned by the DSR-RSPG. Further methodological tools are needed as future work to assess the validity of the composition graph and spline from 5. In contrast, the expression generated by DSR-PQT does not include $d$, and the expression of DSR-VPG excludes both $f$ and $d$. These results indicate that the DSR-PQT and DSR-VPG expressions are unlikely to be valid to accurately model PL from a communication engineering perspective.

### v4

@@ III-B The Indoor and Outdoor Empirical Models
El Chall et al.[8] derived empirical PL models for LoRaWAN (Long-Range Wide Area Network) based on extensive field measurements conducted in Lebanon. Indoor experiments were carried out at Saint Joseph University, while outdoor experiments were performed in Beirut City. The collected data were analyzed using linear regression to derive PL models for both indoor and outdoor environments.
@@
where $n=2.85$ is the PL exponent, $PL_{0}=120.4$ is the reference PL, $b=0.47$ is a fitting parameter, $L_{f}=10$ and $L_{w}=1.41$ are the attenuation factors for floors and walls, respectively.
@@ IV Methodologies
The design decision to adopt KANs and DSR as candidate techniques to enable the automation of PL learning is justified by the shortcomings of typical techniques in learning logarithmic functions [33]. Knowing that existing PL models often involve logarithmic functions in their expression [36], KANs and DSR may have superior performance in approximating such models while being able to produce different flavors of functional interpretability.
Overall, different methods provide varying levels of interpretability. DNNs offer low interpretability through feature importance or activation maps. In contrast, KANs offer a medium level of structural interpretability through their spline-based representations, KAN graphs and edge weights. DSR and KANs (by using Auto-symbolic mapping) can achieve high functional interpretability by expressing models as explicit mathematical functions. Neither level of interpretability implies a direct mapping to individual electromagnetic effects; instead, both focus on making the learned large-scale propagation laws transparent and analyzable and eventually validated by humans as depicted in Figure 1.
@@ IV-A Kolmogorov-Arnold Networks
KANs have recently emerged as an alternative to multi-layer perceptron (MLP) based architectures, demonstrating strong performance on scientific tasks while maintaining interpretability. KANs have strong mathematical foundations based on the Kolmogorov-Arnold representation theorem, which states that any multivariate continuous function can be expressed as a sum of univariate functions [30]. KANs and MLPs implement activation functions in fundamentally different ways: KANs have activation functions on edges, while MLPs have activation functions on nodes. This unique design allows KANs to provide a more transparent and interpretable mapping between inputs and outputs, enhancing explainability while preserving the model’s ability to capture complex patterns.
@@
IV-B Deep Symbolic Regression (DSR)
At the core of DSR [35] are the representation of mathematical expressions as sequences, the development of an autoregressive model to generate these expressions, and the use of a risk-seeking policy gradient approach to train the model to produce more accurate and well-fitting expressions.
@@ IV-B1 Risk-Seeking Policy Gradient (RSPG)
Originally proposed in [35], this approach prioritizes optimizing the best-case reward rather than the average. The loss function is defined as:
@@ IV-B2 Vanilla Policy Gradient (VPG)
This method employs the REINFORCE algorithm [46], where training is performed over the batch $\mathcal{T}$ with the loss function:
@@ IV-B3 Priority Queue Training (PQT)
Proposed by Abolafia [1], this non-RL approach focuses on optimizing best-case performance. It stores samples from each batch in a maximum reward priority queue (MRPQ), and training is conducted on these stored samples using a supervised learning objective:
@@ V Evaluation Methodologies
Figure 2: The flow diagram of developing a PL model using KANs.
@@ V-A1 Synthetic Data
The dataset was generated following the approach described in [3] to approximate the PL models defined by Eq. (1) and Eq. (2) in Section III-A. The parameter ranges, summarized in Table I, were derived from authentic data in studies that comprehensively analyze various scenarios and frequency ranges, highlighting the applicability of $PL^{ABG}$ and $PL^{CI}$ models in various contexts [42, 41, 10]. Specifically, for $PL^{ABG}$ the parameters $\alpha$, $\gamma$, $\beta$, $f$, $d$, and $\chi^{ABG}_{\sigma}$ were used as inputs, while for $PL^{CI}$ $f$, $n$, $d$, and $\chi^{CI}_{\sigma}$ were the inputs. The dataset consists of 1,000 instances. The frequency was converted to hertz for the $PL^{CI}$, following standard practice in the literature.
@@ V-A2 Empirically Measured Data
El Chall et al. [8] conducted measurements in both indoor and outdoor environments. The indoor dataset includes 1,317 instances, and the outdoor dataset contains 787 instances. Input parameters for $PL^{EI}$ included $n_{w}$, $n_{f}$, $d$, and $f$ , while the $PL^{EO}$ used $h_{\mathrm{ED}}$, $d$, and $f$. The corresponding parameter ranges are presented in Table II. For performance evaluation, we compared the automated methods with the models described in Section III-B and developed by the authors in [8] using the real-world measurements [^2] https://zenodo.org/records/1560654.
@@ V-B Training of KANs
We used the original KAN implementation from the pykan library [^3] https://github.com/KindXiaoming/pykan and trained the KANs in a similar way as other ML models, using the 80-20 split of data as depicted in Figure 2. During training, layers and grid sizes are configured to optimize performance. In this process, the splines and weights of the edges connecting these splines, as described in Section IV-A, are learned. Unlike MLPs, KANs are not very deep (i.e., usually about 3 layers with 5 being very deep) and not very wide [30]. Due to their relative shallowness and narrowness, once learned, they can be visualized as a tree structure with the learned activation functions overlayed on the edges and the importance of compositions represented as edge weight. We refer to this structure as tree structure with splines in the workflow illustrated in Figure 2. This alone already gives more intuition on the learned model than in the case of other neural architectures.
As a second step, a mechanism to match the learned splines with actual known functions also exists [30] in the original pykan library. The matching approach relies on a configurable library of elementary functions such as $\sin(\cdot)$, $\exp(\cdot)$, and $\log(\cdot)$. These functions are ranked based on how closely they match the spline using $R^{2}$ and the closest match with the simplest form is preferred. The selected function is then refined using least squares to match the spline as close as possible [30]. This mechanism is implemented in the library’s default auto_symbolic()auto\_symbolic() functionality, which we use and is represented as the Auto-symbolic bloc in Figure 2. After the matching step, the network is retrained with the respective functions by only updating the weights of the edges resulting in the tree structure with functions from Figure 2. This second tree structure can be re-written as a final symbolic expression. As can be seen from the figure, KANs with the auto-symbolic mapping functionality are able to provide a two-level interpretability insight: a tree with spline composition and a symbolic expression. To tune the KANs, we perform a full grid search for each model with $grids\in\{5,8,10,15,20,30,40,50\}$, $steps\in\{50,100,200,300\}$ and $lambda\in\{0.02,0.002,0.0002,0.00002\}$.
@@
(a) Parameter search for ABG.
(b) Parameter search for CI.
Figure 3: KAN model performance as a function of grid size, steps and lambda. The conclusions for indoor and outdoor models are similar to the ABG and CI models.
Hyperparameter selection: To quantify the trade-off between model performance performance and computational complexity, we analyzed the sensitivity of the coefficient of determination ($R^{2}$) to variations in structural complexity, parameterized by the KAN grid size ($G$), steps and $\lambda$ contextualizing the parameters selected in Table III. As illustrated in Figure 3, the performance landscape demonstrates that accuracy does not scale linearly with complexity; rather, it exhibits a rapid saturation characteristic. Specifically, for both the ABG and CI datasets, models with minimal grid densities ($G\in[3,10]$) consistently achieve near-optimal accuracy ($R^{2}>0.98$, indicated by high-intensity markers), comparable to dense models with significantly higher degrees of freedom ($G=50$). This plateau in performance suggests that the underlying physical dynamics can be effectively captured by low-complexity symbolic representations. Considering also the results obtained in our recent work on eCAL, a new metric developed to assess the energy cost of artificial intelligence lifecycle [11], which shows that with the KAN architectures considered in this paper, together with small training data, AutoPL can run on personal laptops. Moreover, for training, it consumes less than 2 Joules of energy on average.
The final selection of tuned hyperparameters for the PL models learned by KANs are summarized in Table III. For both the CI and ABG models, we normalize the training data to maintain consistency with our previous work [3]. In preliminary experiments without normalization, the frequency, expressed in hertz in the CI model, tended to dominate the other input variables, resulting in imbalanced feature contributions.
Figure 4: The flow diagram of developing a PL model using DSR.
@@ V-C Training of DSR
(a) Parameter search for ABG.
(b) Parameter search for CI.
Figure 5: DSR model performance as a function of policy and samples. The conclusions for indoor and outdoor models are similar to the ABG and CI models.
@@
In DSR, a set of constraints can be implemented to narrow the search space, ensuring the generation of meaningful expressions while preserving computational efficiency. The operator library was selected to align with the functional forms commonly found in established PL models (logarithmic, polynomial, and linear operators), while the constraints described below ensure physically meaningful and interpretable outputs. The following constraints were specifically applied. Expression Length Limits: Expressions were constrained to a minimum length of 4 tokens to avoid trivial solutions and a maximum length of 40 tokens to maintain interpretability. Avoiding Redundant Constants: Operators were restricted to prevent all their children from being constants, preventing the generation of expressions that simplify to a single constant. Unary Operator Consistency: Unary operators were restricted from having their inverse as a child (e.g., $\log(\exp(x))$) to prevent the generation of redundant expressions. Trigonometric Operator Composition: Trigonometric operators were restricted from having other trigonometric operators as descendants (e.g., $\sin(1+\cos(x))$) to prevent unnecessary complexity, despite being mathematically valid. Repeat Constraint: This constraint regulates the occurrence of specific tokens (e.g., operators and/or variables) in generated expressions by enforcing minimum and maximum limits. It softly penalizes violations of the minimum threshold to allow exploration while strictly invalidating expressions that exceed the maximum limit. The repeat constraint with $f$ and $d$ was utilized in all cases.
Hyperparameter selection: To quantify the trade-off between model performance and computational complexity, we analyzed the sensitivity of the coefficient of determination ($R^{2}$) to variations in structural complexity, which is shown in Figure 5. This figure is parameterized by the DSR policy and sample size contextualizing the parameters selected in Table IV. In addition, the 3D scatter plots reveal distinct patterns in how the three DSR-based policies perform across varying sample sizes. For the ABG dataset, all three policies start at similar $R^{2}$ values (0.59-0.61) with only 1k samples, but their performance trajectories diverge significantly as sample size increases. RSPG and PQT maintain comparable performance throughout, reaching $R^{2}$ values of 0.92 and 0.93, respectively, at 50k samples, while VPG consistently lags behind, achieving only 0.68 at 5k and 0.88 at 50k samples. This suggests that risk-seeking and priority queue training strategies are more sample-efficient for the ABG model structure. For the CI dataset, an interesting crossover pattern emerges: RSPG achieves the highest $R^{2}$ (0.76) at 1k samples, but PQT surpasses both alternatives at medium sample sizes (0.90 at 5k vs 0.82 for RSPG). At larger sample sizes (20k-50k), all three policies converge to high performance ($R^{2}>0.93$), with RSPG slightly leading at 50k (0.95). The CI dataset appears more amenable to symbolic regression overall, with all policies achieving higher $R^{2}$ values compared to ABG at equivalent sample sizes. Both figures demonstrate that increasing sample size yields diminishing returns beyond 20k-30k samples, where performance begins to plateau across all policies.
Each of the three RNN-based algorithms (RSPG, PQT, and VPG) has one unique algorithm-specific parameter, for which we adopted default values as presented in [35], i.e., the risk factor ($\epsilon=0.05$) for RSPG, the priority queue size ($k=10$) for PQT, and the EWMA coefficient ($\alpha=0.25$) for VPG. In addition to these unique parameters, all three algorithms share a common set of hyperparameters (batch size, learning rate, and entropy weight). We explored various values for these hyperparameters to achieve the best performance. The final tuned hyperparameters are summarized in Table IV.
@@
TABLE V: Performance comparison on ABG model. The interpretability and validity assessment follow the methodology in Section V-D.
[…]
KAN [6,6,1]* 0.39 $\pm$0.74 1.27 $\pm$3.35 2.86 $\pm$3.63 0.99 $\pm$0.00 Tree and splines in Fig. 7 Medium ?
KAN [6,6,1]* Auto-symbolic 3.50 $\pm$1.60 25.19 $\pm$17.70 22.79 $\pm$25.67 0.94 $\pm$0.04 $-291.5\log_{10}{(8.7-3.1\alpha)}+32.2\cos{(1.1\gamma-1.3)}+9\beta-53.3\log_{10}{(9.4-3.9d)}+0.02(0.03f-1)^{2}+313.3-\frac{0.5}{(-f-0.2)^{2}}+8\chi_{\sigma}^{\text{ABG}}$ High
DSR-RSPG 4.57 $\pm$0.25 33.26 $\pm$3.45 11.86 $\pm$0.86 0.92 $\pm$0.01 $20\alpha+10\gamma+\beta+\log_{10}(d)+\frac{\gamma f}{10}+\chi_{\sigma}^{\text{ABG}}$ High
DSR-PQT 5.07 $\pm$0.48 42.09 $\pm$7.06 12.89 $\pm$1.65 0.90 $\pm$0.02 $20\alpha+15\gamma+\beta+\frac{f}{20}+\chi_{\sigma}^{\text{ABG}}$ High
DSR-VPG 7.43 $\pm$1.11 87.38 $\pm$23.52 19.12 $\pm$2.91 0.79 $\pm$0.06 $(\alpha+\gamma)(10+\chi_{\sigma}^{\text{ABG}})$ High
ResNet-MLP 1.68 $\pm$0.24 4.91 $\pm$1.45 15.63 $\pm$21.99 0.99 $\pm$0.00 / none
TabNet 1.61 $\pm$0.35 5.39 $\pm$2.66 12.77 $\pm$15.48 0.99 $\pm$0.00 Feature importance: $\alpha$: 0.3908, $\gamma$: 0.1914, $\beta$: 0.0812, $f$: 0.0614, $d$: 0.2029, $\chi_{\sigma}^{\text{ABG}}$: 0.0723 Low
@@ V-D Evaluation Metrics
To assess the interpretability of the resulting model, we consider the three levels defined in Section IV, namely low, medium, and high. Finally, to assess the validity of the expression, we perform a human expert evaluation, according to the flow depicted in Figure 1. We explicitly examine the final symbolic expressions produced by both DSR and KAN Auto-symbolic to verify whether they preserve the expected monotonic increase of PL with respect to distance and frequency. Expressions containing oscillatory terms (such as sine or cosine of distance or frequency) that violate this fundamental large-scale propagation property are therefore marked as physically invalid in the evaluation tables. Such monotonicity analysis is however, not straightforward or sometimes even possible with other black-box AI models.
V-E DL-based baseline
To establish robust DL baselines for the PL datasets, we implemented two SotA architectures: a ResNet-MLP [20] and TabNet [4]. The ResNet-MLP adapts residual learning for tabular data, featuring an input projection followed by two residual blocks with a hidden dimension of 64, with each block incorporates batch normalization, ReLU activation, and a dropout rate of 0.1 to mitigate overfitting. We also employed TabNet, an attentive interpretable tabular learning architecture, configured with sequential attention. Both models were trained for 150 epochs with a batch size of 32 using the Adam optimizers with learning rates of $10^{-3}$ and $10^{-2}$, respectively. Similar to the proposed methods, to ensure statistical reliability, performance was evaluated using a Monte Carlo cross-validation protocol with 10 independent runs, utilizing an 80/20 train-test split and standard scaling for feature normalization.
@@ VI-A Learning the ABG Model from Synthetic Data
[…]
Among DSR models, DSR-RSPG obtained the best performance (MAE = 4.57 $\pm$0.25, MSE = 33.26 $\pm$3.45, MAPE = 11.86 $\pm$0.86 and $R^{2}$ = 0.92 $\pm$0.01), followed by DSR-PQT (MAE = 5.07 $\pm$0.48, MSE = 42.09 $\pm$7.06, MAPE = 12.89 $\pm$1.65, $R^{2}$ = 0.90 $\pm$0.02) and DSR-VPG (MAE = 7.43 $\pm$1.11, MSE = 87.38 $\pm$23.52, MAPE = 19.12 $\pm$2.91, $R^{2}$ = 0.79 $\pm$0.06). While the DSR models may show improvements at high training samples seizes, the training times are significantly larger, taking even days, compared with KANs where the most complex took hours.
[…]
@@
Looking at the DL baselines, neither produces a symbolic expression. ResNet-MLP remains a complete black box, while TabNet provides only low-level interpretability via feature importance scores ($\alpha\approx 0.39$, $d\approx 0.20$, and $\gamma\approx 0.19$) without revealing the explicit functional relationships governing the PL.
@@
The validity labels reported in Tables V–VIII explicitly account for monotonicity violations with respect to distance and frequency as discussed in Section V-D. We examine the validity of the expressions generated by the KANs and DSR methods. For $PL^{ABG}$, the expressions obtained by KAN [6,6,1] Auto-symbolic and DSR-RSPG are considered valid, as they correctly represent PL as a function of $f$ and $d$. We also notice that the expression learned by the KAN [6,6,1] Auto-symbolic is more complex than the one learned by the DSR-RSPG. Further methodological tools are needed as future work to assess the validity of the composition graph and spline from Figure 7. In contrast, the expression generated by DSR-PQT does not include $d$, and the expression of DSR-VPG excludes both $f$ and $d$. These results indicate that the DSR-PQT and DSR-VPG expressions are unlikely to be valid to accurately model PL from a communication engineering perspective.
TABLE VI: Performance comparison on CI model. The interpretability and validity assessment follow the methodology in Section V-D.
[…]
KAN [4,4,1]* 0.37 $\pm$0.22 1.32 $\pm$1.21 0.29 $\pm$0.16 1.00 $\pm$0.00 Tree and splines in Figure 9 Medium ?
KAN [4,4,1]* Auto-symbolic 3.39 $\pm$2.12 24.55 $\pm$24.41 2.27 $\pm$1.39 0.98 $\pm$0.02 $-280.5\log_{10}{(8.7-1.6n)}+43.8\log_{10}{(1.1d+1.9)}+19.9\log_{10}{(5.2f+9.3)}+50.8\log_{10}{(0.8\chi_{\sigma}^{\text{CI}}+7.5)}+352.2$ High
DSR-RSPG 13.15 $\pm$3.26 296.77 $\pm$137.93 8.63 $\pm$2.05 0.72 $\pm$0.13 $23n+\frac{d}{10}+\log_{10}(f)+40$ High
DSR-PQT 10.40 $\pm$2.69 178.09 $\pm$72.00 6.81 $\pm$1.74 0.83 $\pm$0.07 $23n+\frac{d}{10}+(\chi_{\sigma}^{\text{CI}}+n)\cos(\log_{10}(f))+40$ High
DSR-VPG 14.52 $\pm$2.08 336.70 $\pm$98.19 9.52 $\pm$1.38 0.68 $\pm$0.09 $21n+2\chi_{\sigma}^{\text{CI}}+60$ High
ResNet-MLP 3.59 $\pm$0.71 32.99 $\pm$12.35 2.56 $\pm$0.54 0.97 $\pm$0.01 / None
TabNet 2.21 $\pm$0.41 11.44 $\pm$4.97 1.54 $\pm$0.31 0.99 $\pm$0.00 Feature importance: $f$: 0.2891, $n$: 0.3825, $d$: 0.2136, $\chi_{\sigma}^{\text{CI}}$: 0.1147 Low